A LED company claims that the average life of the LED light bulbs it

manufactures is 1, 500 hours with a standard deviation of 500 hours. If a random sample of 40 bulbs is chosen, what is the probability that the sample mean will be:


1
Expert's answer
2021-06-03T06:57:27-0400

Let "X=" the sample mean: "X\sim N(\mu, \sigma^2/n)." Then "Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0, 1)"

Given "\mu=1500\ h, \sigma=500\ h, n=40."


a. greater than 1, 400 hours?


"P(X>1400)=1-P(X\leq 1400)"

"=1-P(Z\leq \dfrac{1400-1500}{500/\sqrt{40}})"

"\approx1-P(Z\leq-1.264911)\approx0.897048"

b. less than 1, 400 hours?


"P(X<1400)=P(Z< \dfrac{1400-1500}{500/\sqrt{40}})"

"\approx P(Z<-1.264911)\approx0.102952"

Check


"0.897048+0.102952=1"